OpenStudy (nincompoop):

proof:

OpenStudy (nincompoop):

if a < b and c < 0, then ac > bc proof: since c < 0 0 = c + (-c) < 0 + (-c) = -c so -c is positive, hence by (x), we get -ac < -bc, -ac + (bc + ac) < -bc + (bc + ac) hence bc < ac, and this is equivalent to ac > bc I think there's some application of the following axioms: transitive law: if a < b and b < c, then a < c law of trichotomy: for every real number a, one and only of the following alternatives holds: a = 0, or a < 0, or 0 < a; if a < b, then a + c < b + c; if a < b and 0 < c, then ac < bc

OpenStudy (bahrom7893):

Did you get that from my notes?

OpenStudy (bahrom7893):

I recall doing something similar.

OpenStudy (nincompoop):

I don't get this part: 0 = c + (-c) < 0 + (-c) = -c No, it's from a review of real numbers ineq. and absolute values.

OpenStudy (bahrom7893):

Oh okay.. Let's try to break it up: 0 = c + (-c) right?

OpenStudy (nincompoop):

yeah

OpenStudy (bahrom7893):

So we can simplify it down to: 0 < 0 + (-c), because: We are given that c < 0, meaning it's negative. Therefore -c > 0 (double negative is a positive), so this line is simply telling you that: 0 < -c, or -c > 0

OpenStudy (bahrom7893):

Did you follow?

OpenStudy (bahrom7893):

If not I think I can make it even easier.

OpenStudy (nincompoop):

yeah by assuming that -c is positive, but that would mean we're blowing away the fact that -1 (-c) = c > 0

OpenStudy (nincompoop):

OHHHH I GET IT NOW

OpenStudy (bahrom7893):

That Eureka moment in math... ahhhh..

OpenStudy (nincompoop):

ya! thanks. that -c is an insertion or expansion using the existence of subtraction axiom. -a such that a + (-a) = 0